OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x^3 + 4 and g(x) = cuberoot(x-4)

OpenStudy (anonymous):

i know how to tell if they are inverses but I don't know how to show it the way they want me to.

OpenStudy (anonymous):

f(g(x))=((x-4)^(1/3))^3+4=x-4+4=x g(f(x))=(x^3+4-4)^(1/3)=(x^3)^(1/3)=x so f and g are inverses.

OpenStudy (anonymous):

Ohhhhhh wow thanks! :)